40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Nov 6 07:26:54 2010, yuta, Summary, IOO, reduced common mode displacement of the beam through MC1 to MC3 A2LMCalign.png
    Reply  Tue Nov 9 05:40:12 2010, yuta, Summary, IOO, MC aligning going on MCalignNov9.png
       Reply  Wed Nov 10 02:51:35 2010, yuta, Summary, IOO, limitation of current MC aligning MCalignNov9.png
          Reply  Wed Nov 10 11:46:19 2010, Koji, Summary, IOO, limitation of current MC aligning 
             Reply  Wed Nov 10 12:21:18 2010, yuta, Summary, IOO, limitation of current MC aligning 
                Reply  Wed Nov 10 14:28:33 2010, Koji, Summary, IOO, limitation of current MC aligning 
Message ID: 3884     Entry time: Wed Nov 10 02:51:35 2010     In reply to: 3883     Reply to this: 3885
Author: yuta 
Type: Summary 
Category: IOO 
Subject: limitation of current MC aligning 

(Suresh, Yuta)

Summary:
  We need MC to be locked and aligned well to align other in-vac optics.
  We continued to align the incident beam so that the beam passes the actuation nodes of MC1 and MC3.
  From the previous measurement, we found that beam height at IM1 has to be increased by ~3cm.
  Today, we increased it by ~1cm and achieved about 1/3 of the required correction.
  But we cannot proceed doing this because the beam is hitting IM1 at the edge already.

What is the goal of this alignment?:
  If the beam doesn't hit MC optics in the center, we see angle to length coupling, which is not good for the whole interferometer.
 
  Also, if the beam is tilted so much, transmitted beam though MC3 cannot go into FI at right after MC3.
  Say, FI has an aparture of 3mm and MC3-FT distance is 300mm. The beam tilt should be smaller than 3/300 rad. MC1-MC3 distance is 200mm, so the displacement at each mirror should be smaller than ~1mm.
  1mm is about 7% (see Koji's elog #2863) TO_COIL gain imbalance in A2L measurement.
 
  We are currently assuming that each coils are identical. If they have 5% variance, it is meaningless to try to reduce the beam displacement less than ~5%.

  So, we set the goal to 7%.

What we did:

  1. Leveled the MC table.

  2. Measured the table height using DISTO D3 laser gauge.
    PSL table 0.83m (+-0.01m)
    OMC table 0.82m
    MC table  0.81m

  3. Using the last steering mirror(SM@PSL) and IM1, tilted the beam vertically

Result:

MCalignNov9.png

  At t=0 (this morning), the beam tilt was ~40%/(MC1-MC3 distance). Now, it is ~30%/(MC1-MC3 distance).
  30%/(MC1-MC3 distance) is ~5/200 rad.

Plan:

 We have to somehow come up with the next story. Too much vertical tilt. What is wrong? Table leveling seems OK.
 - measure in-vac beam height
 - maybe OSEMs are badly aligned. we have to check that.

ELOG V3.1.3-